comsumer consuming two goods x and y has utility function v=x➋+y➌.If the price of the two commodities 6 and 2 respectively and his budget is birr .40 find the quantites x and y which will maximize utility


1
Expert's answer
2021-07-11T22:32:02-0400

As the price of y is less ,more number of y can be bought.In the utility function, part of y has a greater power.This means it is better to buy y with the whole budget to maximize utility.

"X=0 \\and \\y=\frac{40}{2}=20"


"MRS=\frac{MUx}{MUy}=\frac{2x}{2y}=0 ,(x=0)"